# M2SVid: End-to-End Inpainting and Refinement for Monocular-to-Stereo Video Conversion

### v1

@@ Abstract
We tackle the problem of monocular-to-stereo video conversion and propose a novel architecture for inpainting and refinement of the warped right view obtained by depth-based reprojection of the input left view. We extend the Stable Video Diffusion (SVD) model to utilize the input left video, the warped right video, and the disocclusion masks as conditioning input to generate a high-quality right camera view. In order to effectively exploit information from neighboring frames for inpainting, we modify the attention layers in SVD to compute full attention for discoccluded pixels. Our model is trained to generate the right view video in an end-to-end manner by minimizing image space losses to ensure high-quality generation. Our approach outperforms previous state-of-the-art methods, obtaining an average rank of 1.43 among the 4 compared methods in a user study, while being 6×6$\times$6 × faster than the second placed method.
@@ 1 Introduction
Emerging technologies such as Mixed-Reality headsets and glasses allows users to easily experience immersive content, thanks to the use of separate displays for left and right eyes. Rendering a video from left and right eye viewpoints on corresponding left and right displays can create a stereoscopic 3D effect, giving viewers a perceived sense of depth. However, capturing such stereoscopic videos usually requires specialized cameras that can capture both left and right eye perspectives simultaneously. While manual monocular-to-stereo conversion is possible, it is a costly and time-consuming process that requires significant manual effort. Therefore, there is a need for automated, fast and high-quality monocular-to-stereoscopic video conversion methods that can enable large scale conversion of videos.
@@
Recent advancements in video generation [5, 4] as well as monocular depth estimation [83, 31] have led to significant improvements in the monocular-to-stereo conversion quality. A common approach is to use depth maps generated by a monocular depth model to warp the input video to a virtual right camera and then inpaint the disoccluded regions [49, 10, 68]. SVG [10] repurposes a pre-trained video diffusion model for the inpainting task, while  [49] trains a custom model. Rather than only modifying discoccluded regions, the recent concurrent work StereoCrafter [68] finetunes a video diffusion model [3] to inpaint and refine the full warped video, using only the warped view and discocclusion masks as input. While obtaining improved results, StereoCrafter still leverages an architecture primarily designed for monocular video inpainting task, without exploiting the constraints specific to stereoscopic video-refinement. Moreover, the inference requires multiple diffusion steps (default 25) leading to high computational cost.
@@
In this work, we propose a novel architecture for efficient inpainting and refinement in the monocular-to-stereo conversion task. Our approach is designed to leverage important additional cues available in this task. Firstly, in the areas that are not disoccluded, the right view largely resembles the left one, where the image content is just horizontally shifted. Thus, even if depth-based warping introduces artifacts in the generated right view, we conjecture that most of them can be fixed by relying on left view information. Secondly, the disoccluded regions are generally small in most cases. Furthermore in the presence of camera or scene motion, the disoccluded regions in one frame are often visible at nearby frames, thus simplifying the inpainting task into a spatial-temporal matching problem. The aforementioned aspects motivate us to design M2SVid, a specialized architecture which performs inpainting and refinement in an feed-forward manner, without any iterative diffusion steps.
@@
Our main contributions are the following. First, we extend the Stable Video Diffusion (SVD) architecture by employing the input left view, in addition to the warped right view and disocclusion mask as conditioning for the refinement task. This makes it easier for the model to propagate information from the uncorrupted left view to the target right view, thus preserving high frequency details such as fine structures. Secondly, rather than performing temporal attention only over the same spatial locations as in SVD [3], we compute full cross-attention for disoccluded pixels. This provides greater flexibility to propagate useful information from either nearby temporal frames or the associated left view, with a limited computational overhead. Finally we train our model in an end-to-end manner using perceptual and fidelity based losses, enabling high quality inpainting and refinement (See Fig. 1).
@@
We perform quantitative evaluation, qualitative analysis, as well as a human perception user study to evaluate our method. It shows that our M2SVid outperforms previous state-of-the-art StereoCrafter and SVG, obtaining an average rank of 1.43 / 4.0 in the user study, while being 6×6$\times$6 × and 600×600$\times$600 × faster than StereoCrafter and SVG respectively.
@@ 2.1 Novel View Synthesis
Early methods for novel-view synthetis performed 3D reconstruction of the full scene given dense input views [65, 28, 84], which can then be used to render novel views. While Multi-Plane Image (MPI)-based approaches [74, 18, 73] were successfully used to generate multiplane representations from single images, NeRFs [52] marked a milestone in the field, and spawned many efforts to improve quality [1, 2], reduce latency [54, 19, 78, 85, 55], extend to large-scale scenes [75], explicit representation [32], pose-free setup [40, 77, 8, 51, 81, 72], or fewer inputs [23, 33, 56, 11, 59, 72, 80], including just a single image [50, 42, 43, 44]. Approaches such as CAT3D [15], NVS-Diffusion [29], MultiDiff [53] directly generate the queried camera view using a diffusion models. A few works also extend these to videos [79]. While obtaining promising results, the generated views can still have artifacts due to the inherent difficulties of rendering views with large viewpoint changes.
@@ 2.2 Monocular-to-Stereo Conversion
Compared to the general novel-view synthesis problem, Monocular-to-Stereo conversion aims to only render a fixed right camera view given an input left view. This can allow viewing any monocular content in 3D on mixed reality headsets, enabling a more immersive experience. The majority of previous works can be categorized into two families. One line of work aims to directly generate the right image/video given the left input [82, 82, 67, 36]. A notable example is Deep3D [82], which leverages an internal disparity representation to directly render the right video given the left one. The second type of approaches employ an explicit depth map to reproject the left image to the right camera, and then inpaint the disoccluded regions [69, 24, 49, 10, 68, 35, 76, 10]. Notably, Wang et al. [76] learn a diffusion model for inpainting disoccluded pixels using self-supervision with random cycle rendering. while SVG [10] uses a pre-trained video diffusion model for inpainting. While the method can achieve spatio-temporal consistency in the output, the inpainting result can be incorrect due to the lack of task-specific fine tuning. Mehl et al. [49] aim to instead perform a geometry aware inpainting of the disoccluded areas using local background information, rather than generating arbitrarily realistic content. StereoCrafter [68], which is concurrent to our work, fine-tunes a video diffusion model for inpainting the disoccluded areas, on a large stereo dataset (not released). Another concurrent work SpatialDreamer [47] mitigates the necessity of paired stereoscopic training data by proposing a self-supervised training framework using a forward-backward rendering mechanism. Note that a number of the above methods do not release code or model, and are often trained on private collected data, making a fair comparison difficult.
@@ 2.3 Diffusion Models
Diffusion Models [20, 70, 60] are a family of generative models that iteratively denoise an image to generate the output. They have been tremendously successful for the text-to-image [20, 60, 63, 87, 58] and text-to-video [4, 21] tasks. They have also been successfully employed for a wide variety of computer vision tasks, thanks to the strong image priors learned by the models. These include image generation applications such as image [46, 9, 41] and video [92, 7, 37] inpainting, novel view synthesis [53, 27], as well as other pixel-wise prediction tasks such as monocular depth estimation [31, 22, 16, 66], semantic segmentation [30], or normal estimation [14]. Conventionally, diffusion models employ multiple denoising steps during inference, leading to large computations cost. A number of approaches aim to train one-step models to address this, using knowledge distillation [86, 90], GAN training [48], low-rank adaptors [90]. Garcia et al. [16] instead finetune a pre-trained diffusion model to perform direct feed-forward monocular depth estimation using end-to-end training. In this work, we show that such an end-to-end training strategy can also be employed for inpainting disoccluded regions in monocular-to-stereo video conversion task.
@@ 3 Conditional Latent Diffusion Models
Here, we provide a brief background on diffusion models employed in our method. Denoising Diffusion Probabilistic Models (DDPMs) [20] are generative models trained to map a simple noise distribution $p_{T}$ to the data distribution $p_{0}$, by reversing a stochastic forward process $p_{t}$, $t=1,...,T$. A denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t)$ is trained to generate an image from pure noise by progressively denoising inputs $x_{t}$ at time stamp $t$. In order to reduce computational complexity, Latent Diffusion Models (LDMs) [60] operate in a latent space of a Variational Autoencoder (VAE) [34]. Conditional diffusion models [62, 88] extend the denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t,c)$ to be conditioned on additional input $c$, such as text [62], images [62], etc. to control the generation process.
Training and inference with conditional LDMs:  During training, given a data sample $\mathbf{x}$ (e.g., an image or video) and its corresponding conditioning input $c$ (e.g., text or another image), the data sample $\mathbf{x}$ is first encoded into a latent representation $\mathbf{z}=E(\mathbf{x})$ using VAE encoder $E$. The latent representation $\mathbf{z}$ is perturbed through a forward diffusion process: $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ is Gaussian noise, and $\bar{\alpha}_{t}$ controls the noise schedule. Instead of predicting the noise $\boldsymbol{\epsilon}$ directly, diffusion models usually leverage v-parameterization [64], where the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{z}_{t},t,c)$ learns to predict: $\boldsymbol{\mathbf{v}}=\alpha_{t}\boldsymbol{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}.$ Therefore the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}$, typically a U-Net [61], is trained to reconstruct $\boldsymbol{\mathbf{v}}$, by optimizing the objective:
@@
Diffusion as feed-forward models:  Garcia et al. [16] propose to use pre-trained diffusion U-Net model as a feed-forward models for pixel-to-pixel tasks such as depth and normal estimation. In this scenario, the timestep $t$ is not sampled anymore and fixed as $t=T$ to train the model for single-step prediction. The noise $\boldsymbol{\epsilon}$ is additionally replaced by the mean of the noise distribution, i.e., zero. With $t=T$, we get $\bar{\alpha}_{T}\approx 0$, and thus $\mathbf{z}_{T}=\sqrt{\bar{\alpha}_{T}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{T}}\boldsymbol{\epsilon}\approx\mathbf{0}$ and $\boldsymbol{\mathbf{v}}=\alpha_{t}\boldsymbol{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}\approx-\mathbf{z}.$ As such, the model directly learns to predict the clean output $\boldsymbol{\mathbf{v}}\approx-\mathbf{z}$ starting from zero noise vector $\mathbf{z}_{T}=\mathbf{0}$. With such timestep sampling and assumptions, the standard loss (Eq. 1) converges to the $L_{2}$ loss in the latent space between the ground truth $-\mathbf{z}$ and the predicted latents $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{0},T,c)$:
@@
However, [16] proposes decoding the predicted latents, $\mathbf{\hat{z}}=-\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{0},T,c)$, using a VAE decoder $D$ to obtain the reconstructed output $\mathbf{\hat{x}}=D(\mathbf{\hat{z}})$. Instead of minimizing the latent space error Eq. 2, they propose to optimize a task-specific loss directly in image space with $\mathcal{L}=L_{task}(\mathbf{x},\mathbf{\hat{x}})$.
@@ 4 Monocular-to-Stereo Conversion Pipeline
In this work, we tackle the problem of converting any 2D video to a 3D stereoscopic video. Given a monocular video $V\in\mathbb{R}^{N\times H\times W\times 3}$, containing N frames of resolution $H\times W$, the goal is to transform it into a stereoscopic video pair $(V_{l},V_{r})$, where $V_{l},V_{r}\in\mathbb{R}^{N\times H\times W\times 3}$ represent the views corresponding to the left and right eyes, respectively. Following prior work, we assume that the input video $V$ corresponds to the left-eye view, i.e., $V_{l}=V$, and therefore, the goal is to only generate the right-eye view $V_{r}$ from $V$. Note that multiple possible right-eye views $V_{r}$ exist depending on the baseline, i.e. the distance between cameras. Hence we condition the generation on a maximum desired disparity $D_{\text{max}}$ between $V_{l}$ and $V_{r}$. This can be provided by the user to control the amount of stereoscopic effect in practical applications.
An overview of the pipeline is shown in Fig. 2. We follow the commonly employed strategy [49, 68] of performing depth-based reprojection to generate an estimate of the right view, as described next.
@@
Video inpainting and refinement:  The final step in our pipeline is to generate a high quality right view given the warped video $V_{r}^{warp}$. As mentioned before, the warped right video $V_{r}^{warp}$ contains holes introduced by disocclusions. Furthermore, the warped right video $V_{r}^{warp}$ can also contain artifacts introduced due to reprojection and depth errors. We use a Video Diffusion Model to inpaint the missing regions, as well as correct the artifacts present in the warped right views.
In this work, we mainly focus on the final “video inpainting and refinement” task. To this end, we introduce an efficient end-to-end model to inpaint the disoccluded regions and fix reprojection artifacts. Our architecture is described in detail in the next section.

### v2

@@ Abstract
We tackle the problem of monocular-to-stereo video conversion and propose a novel architecture for inpainting and refinement of the warped right view obtained by depth-based reprojection of the input left view. We extend the Stable Video Diffusion (SVD) model to utilize the input left video, the warped right video, and the disocclusion masks as conditioning input to generate a high-quality right camera view. In order to effectively exploit information from neighboring frames for inpainting, we modify the attention layers in SVD to compute full attention for discoccluded pixels. Our model is trained to generate the right view video in an end-to-end manner without iterative diffusion steps by minimizing image space losses to ensure high-quality generation. Our approach outperforms previous state-of-the-art methods, being ranked best $2.6\times$ more often than the second-place method in a user study, while being $6\times$ faster. [^1] To be published at 3DV 2026. When citing this work, please refer to the final version published in IEEE Xplore. Cite as: Nina Shvetsova, Goutam Bhat, Prune Truong, Hilde Kuehne, Federico Tombari. “M2SVid: End-to-End Inpainting and Refinement for Monocular-to-Stereo Video Conversion”. In International Conference on 3D Vision, IEEE, 2026.
@@ 1 Introduction
Emerging technologies such as Mixed-Reality headsets and glasses allows users to easily experience immersive content, thanks to the use of separate displays for left and right eyes. Rendering videos from left and right eye viewpoints on these displays creates a stereoscopic 3D effect, giving viewers a sense of depth. However, capturing such stereoscopic videos usually requires specialized cameras that can capture both left and right eye perspectives simultaneously. While manual monocular-to-stereo video conversion is possible, it is costly and time-consuming. This drives the need for automated, fast, and high-quality conversion methods that can enable large scale conversion of videos.
@@
Recent advancements in video generation [chen2024videocrafter2, blattmann2023stable] as well as monocular depth estimation [depth_anything_v2, ke2024repurposing] have led to significant improvements in the monocular-to-stereo conversion quality. A common approach is to use depth maps generated by a monocular depth model to warp the input video to a virtual right camera and then inpaint the disoccluded regions [mehl2024stereo, dai2024svg, shi2024stereocrafter]. SVG [dai2024svg] repurposes a pre-trained video diffusion model for the inpainting task, while  [mehl2024stereo] trains a custom model. Rather than only modifying discoccluded regions, the recent concurrent work StereoCrafter [shi2024stereocrafter] finetunes a video diffusion model [Blattmann2023StableVD] to inpaint and refine the full warped video, using only the warped view and discocclusion masks as input. While obtaining improved results, StereoCrafter still leverages an architecture primarily designed for monocular video inpainting task, without exploiting the constraints specific to stereoscopic video-refinement. Moreover, the inference requires many diffusion steps leading to high computational cost.
@@
In this work, we propose a novel architecture for efficient inpainting and refinement in the monocular-to-stereo conversion task. Our approach is designed to leverage important additional cues available in this task. Firstly, in the areas that are not disoccluded, the right view largely resembles the left one, where the image content is just horizontally shifted. Thus, even if depth-based warping introduces artifacts in the generated right view, we conjecture that most of them can be fixed by relying on left view information. Secondly, the disoccluded regions are generally small in most cases. Furthermore in the presence of camera or scene motion, the disoccluded regions in one frame are often visible at nearby frames, thus simplifying the inpainting task into a spatial-temporal matching problem. The aforementioned aspects motivate us to design M2SVid (Monocular-to-Stereo Video conversion), a specialized architecture which performs inpainting and refinement in an feed-forward manner, without any iterative diffusion steps.
@@
Our main contributions are the following. First, we extend the Stable Video Diffusion (SVD) architecture by employing the input left view, in addition to the warped right view and disocclusion mask as conditioning for the refinement task. This makes it easier for the model to propagate information from the uncorrupted left view to the target right view, thus preserving high frequency details such as fine structures. Secondly, rather than performing temporal attention only over the same spatial locations as in SVD [Blattmann2023StableVD], we compute full cross-attention for disoccluded pixels. This provides greater flexibility to propagate useful information from either nearby temporal frames or the associated left view, with a limited computational overhead. Thirdly, we train our model in an end-to-end manner using perceptual and fidelity based losses, enabling high quality inpainting and refinement (See Fig. 1). Finally, unlike other methods, we train on publicly available datasets, making our approach reproducible and enabling fairer comparisons.
@@
We perform quantitative evaluation, qualitative analysis, as well as a human perception user studies to evaluate our method. It shows that our M2SVid outperforms previous state-of-the-art StereoCrafter and SVG, being ranked best $2.6\times$ more often than StereoCrafter and $4.75\times$ more often than SVG, while running $6\times$ and $600\times$ faster, respectively.
@@ 2.1 Novel View Synthesis
Early methods for novel-view synthetis performed 3D reconstruction of the full scene given dense input views [colmap, KarHM17, YaoLLFQ18], which can then be used to render novel views. While Multi-Plane Image (MPI)-based approaches [tucker2020single, han2022single, single_view_mpi] were successfully used to generate multiplane representations from single images, NeRFs [mildenhall2020nerf] marked a milestone in the field, and spawned many efforts to improve quality [mipnerf, mipnerf360], reduce latency [mueller2022instant, HedmanSMBD21, Nex, PlenOctrees, NeffSPKMCKS21], extend to large-scale scenes [Turki_2022_CVPR], explicit representation [kerbl20233d], pose-free setup [barf, nerfmm, GARF, GNerF, sinerf, sparf], or fewer inputs [dietnerf, infonerf, Regnerf, DSNerf, densedepth, sparf, reconfusion], including just a single image [Melas-KyriaziL023, LiuXJCTXS23, LiuSCZXW00G024, LiuWHTZV23]. Approaches such as CAT3D [gao2024cat3d], NVS-Diffusion [Karras2024edm2], MultiDiff [Muller_2024_CVPR] directly generate the queried camera view using a diffusion models. A few works also extend these to videos [gao2024cat4d]. While obtaining promising results, the generated views can still have artifacts due to the inherent difficulties of rendering views with large viewpoint changes.
@@ 2.2 Monocular-to-Stereo Conversion
Unlike general novel-view synthesis, monocular-to-stereo conversion renders a fixed right view from a left view, enabling monocular content to be experienced in 3D on mixed-reality headsets. Most prior works can be grouped into two families. One line of work aims to directly generate the right image/video given the left input [xie2016deep3d, xie2016deep3d, shi2024immersepro, Lee2017Automatic2C]. A notable example is Deep3D [xie2016deep3d], which leverages an internal disparity representation to directly render the right video given the left one. The second type of approaches employ an explicit depth map to reproject the left image to the right camera, and then inpaint the disoccluded regions [Shih3DP20, Jampani2021SLIDESI, mehl2024stereo, dai2024svg, shi2024stereocrafter, konrad2013learning, wang2023learning, dai2024svg]. Notably, Wang et al[wang2023learning] learn a diffusion model for inpainting disoccluded pixels using self-supervision with random cycle rendering. SVG [dai2024svg] uses a pre-trained video diffusion model for inpainting. While the method can achieve spatio-temporal consistency in the output, the inpainting result can be incorrect due to the lack of task-specific fine-tuning. Mehl et al[mehl2024stereo] aim to instead perform a geometry aware inpainting of the disoccluded areas using local background information, rather than generating arbitrarily realistic content. StereoCrafter [shi2024stereocrafter], which is concurrent to our work, fine-tunes a video diffusion model for inpainting the disoccluded areas, on a large stereo dataset (not released). Another recent work SpatialDreamer [spatialdreamer] mitigates the necessity of paired stereoscopic training data by proposing a self-supervised training framework using a forward-backward rendering mechanism. Note that a number of the above methods do not release code or model, and are often trained on private collected data, making a fair comparison difficult.
@@ 2.3 Diffusion Models
Diffusion Models [ho2020denoising, song2020denoising, rombach2022high] are generative models that iteratively denoise an input to produce the output. They have been tremendously successful for the text-to-image [ho2020denoising, rombach2022high, Saharia2022PhotorealisticTD, Zhang2023AddingCC, Ramesh2022HierarchicalTI] and text-to-video [blattmann2023stable, Ho2022ImagenVH] tasks. They have also been successfully employed for diverse computer vision tasks, thanks to the strong image priors learned by the models. These include image [Lugmayr2022RePaintIU, Corneanu2024LatentPaintII, Liu2023ImageIV] and video [zhang2023avid, cherel2023infusion, lee2025videoinpainter] inpainting, novel view synthesis [Muller_2024_CVPR, kalischek2025cubediffrepurposingdiffusionbasedimage], as well as monocular depth estimation [ke2024repurposing, hu2024depthcrafter, garcia2024fine, shao2024learningtemporallyconsistentvideo], semantic segmentation [Kawano2024MaskDiffusionEP], or normal estimation [fu2024geowizard]. Conventionally, diffusion models employ multiple denoising steps during inference, leading to large computations cost. A number of approaches aim to train one-step models to address this, using knowledge distillation [song2024sdxs, Zhang2023HiPAEO], GAN training [Mao2024OSVOS], low-rank adaptors [Zhang2023HiPAEO]. Garcia et al[garcia2024fine] instead finetune a pre-trained diffusion model to perform direct feed-forward monocular depth estimation using end-to-end training. In this work, we show that such an end-to-end training strategy can also be employed for inpainting disoccluded regions in monocular-to-stereo video conversion task.
@@ 3 Conditional Latent Diffusion Models
Here, we provide a brief background on diffusion models employed in our method. Denoising Diffusion Probabilistic Models (DDPMs) [ho2020denoising] are generative models trained to map a simple noise distribution $p_{T}$ to the data distribution $p_{0}$, by reversing a stochastic forward process $p_{t}$, t=1,,Tt=1,...,T. A denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t)$ is trained to generate an image from pure noise by progressively denoising inputs $x_{t}$ at time stamp $t$. In order to reduce computational complexity, Latent Diffusion Models (LDMs) [rombach2022high] operate in a latent space of a Variational Autoencoder (VAE) [kingma2013auto]. Conditional diffusion models [saharia2022photorealistic, zhang2023adding] extend the denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t,c)$ to be conditioned on additional input $c$, such as text [saharia2022photorealistic], images [saharia2022photorealistic], etc. to control the generation process.
Training and inference with conditional LDMs:  During training, given a data sample $\mathbf{x}$ (e.g., an image or video) and its corresponding conditioning input $c$ (e.g., text or another image), the data sample $\mathbf{x}$ is first encoded into a latent representation $\mathbf{z}=E(\mathbf{x})$ using VAE encoder $E$. The latent representation $\mathbf{z}$ is perturbed through a forward diffusion process: $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ is Gaussian noise, and $\bar{\alpha}_{t}$ controls the noise schedule. Instead of predicting the noise $\boldsymbol{\epsilon}$ directly, diffusion models usually leverage v-parameterization [salimans2022progressive], where the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{z}_{t},t,c)$ learns to predict: $\boldsymbol{\mathbf{v}}=\alpha_{t}\boldsymbol{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}.$ Therefore the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}$, typically a U-Net [ronneberger2015u], is trained to reconstruct $\boldsymbol{\mathbf{v}}$, by optimizing the objective:
@@
Diffusion as feed-forward models:  Garcia et al[garcia2024fine] propose to use pre-trained diffusion U-Net model as a feed-forward models for pixel-to-pixel tasks such as depth and normal estimation. In this scenario, the timestep $t$ is not sampled anymore and fixed as $t=T$ to train the model for single-step prediction. The noise $\boldsymbol{\epsilon}$ is additionally replaced by the mean of the noise distribution, i.e., zero. With $t=T$, we get $\bar{\alpha}_{T}\approx 0$, and thus $\mathbf{z}_{T}=\sqrt{\bar{\alpha}_{T}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{T}}\boldsymbol{\epsilon}\approx\mathbf{0}$ and $\boldsymbol{\mathbf{v}}=\alpha_{t}\boldsymbol{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}\approx-\mathbf{z}.$ As such, the model directly learns to predict the clean output $\boldsymbol{\mathbf{v}}\approx-\mathbf{z}$ starting from zero noise vector $\mathbf{z}_{T}=\mathbf{0}$. With such timestep sampling and assumptions, the standard loss (Eq. 1) converges to the $L_{2}$ loss in the latent space between the ground truth $\mathbf{z}$ and the predicted latents $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{0},T,c)$:
@@
[…]
@@ 4 Monocular-to-Stereo Conversion Pipeline
In this work, we tackle the problem of converting any 2D video to a 3D stereoscopic video. Given a monocular video $V\in\mathbb{R}^{N\times H\times W\times 3}$, containing N frames of resolution $H\times W$, the goal is to transform it into a stereoscopic video pair $(V_{l},V_{r})$, where $V_{l},V_{r}\in\mathbb{R}^{N\times H\times W\times 3}$ represent the views corresponding to the left and right eyes, respectively. Following prior work, we assume that the input video $V$ corresponds to the left-eye view, i.e., $V_{l}=V$, and therefore, the goal is to only generate the right-eye view $V_{r}$ from $V$. Note that multiple possible right-eye views $V_{r}$ exist depending on the baseline, i.e. the distance between cameras. Hence we condition the generation on a maximum desired disparity $D_{\text{max}}$ between $V_{l}$ and $V_{r}$, that can be set by the user to control the stereoscopic effect in practical applications.
An overview of the pipeline is shown in Fig. 2. We follow the commonly employed strategy [mehl2024stereo, shi2024stereocrafter] of performing depth-based reprojection to generate an estimate of the right view, as described next.
@@
Video inpainting and refinement:  The final step in our pipeline is to generate a high quality right view given the warped video $V_{r}^{warp}$. As mentioned before, the warped right video $V_{r}^{warp}$ contains holes introduced by disocclusions. Furthermore, the warped right video $V_{r}^{warp}$ can also contain artifacts introduced due to reprojection and depth errors. In this work, we mainly focus on this final “video inpainting and refinement” task. To this end, we introduce an efficient end-to-end model to inpaint the disoccluded regions and fix reprojection artifacts. Our architecture is described in detail in the next section.